The common stock of a firm paid a dividend of $1.75 in the year just ended. Suppose their dividend is expected to grow at a rate of 10% in the coming year, 8% in year two, and at a rate of 4% annually thereafter. If the required rate of return is 10%, what is the current value of their stock?
Answer #1

D1 = D0 x (1 + g) = 1.75 x (1 + 10%) = 1.925, D2 = D1 x (1 + g) = 1.925 x (1 + 8%) = 2.08,

Price in year 3, P2 = D2 x (1 + g) / (r - g) = 2.08 x (1 + 4%) / (10% - 4%) = $36.04

Current Value today, P0 = D1 / (1 + r) + (D2 + P2) / (1 + r)^2

= 1.925 / 1.10 + (2.08 + 36.04) / 1.10^2

= $33.25